OpenStudy (needhelpp101):

You make the following measurements of an object – 22 kg and 42 m3. What would the object’s density be? Show your work for credit and include final units.

OpenStudy (photon336):

\[Density = \frac{ mass }{ volume }\]

OpenStudy (photon336):

The density can be measured in; Mass = kilograms Volume = meters^3 there are other units used for density.

OpenStudy (photon336):

\[\frac{ Kilograms}{ Meters ^{3} }\]

OpenStudy (photon336):

@Needhelpp101 does this make sense?

OpenStudy (needhelpp101):

42/22 = 1.909 QED

OpenStudy (photon336):

22 = kilograms and that's your mass 42 = your volume in meters cubed. make sure you look at the formula b.c sometimes if you rush in you'll get the wrong answer. the formula says density = mass/volume so 22kg/42m^3 = your density

OpenStudy (needhelpp101):

k